OpenStudy (anonymous):

Multiply. Express your answer in lowest terms. -5/6*(-3/10)

OpenStudy (anonymous):

@Susanna

OpenStudy (anonymous):

@mathmale

Mehek (mehek14):

\(\dfrac{-5}{6}*\dfrac{-3}{10}\) ?

OpenStudy (anonymous):

i think the answer is 1/4

Mehek (mehek14):

15/60 = 1/4 correct :)

OpenStudy (anonymous):

ok thxs

Mehek (mehek14):

yw :)